OpenStudy (anonymous):

Recently , a new-mint copy of the Baltimore Orioles 1966 Yearbook was auctioned for $15. The yearbook cost $.50 in 1966. Write a ratio of the auction price to the original price. Find the percent.

OpenStudy (anonymous):

Would it just be .50/15.00= 0.033x100= 3.33 % ???

OpenStudy (anonymous):

Calculation wrong.....50/15 != 0.033

OpenStudy (anonymous):

So it would still be 3.33 % is that correct ?

OpenStudy (anonymous):

Nooo. Ratio is 15 / 0.50 = 30 and this is 3,000%

OpenStudy (anonymous):

So the auction price to the original price is 3,000 % ? that doesn't sound right to me.

OpenStudy (anonymous):

maybe 30% ?